# Microbenchmarking NVIDIA’s Blackwell Architecture: An in-depth Architectural Analysis

### v1

@@ Abstract
As GPU architectures rapidly evolve to meet the overcoming demands of exascale computing and machine learning, the performance implications of architectural innovations remain poorly understood across diverse workloads. NVIDIA’s Blackwell (B200) generation introduce significant architectural advances including the 5th generation tensor cores, tensor memory (TMEM), decompression engine (DE), and dual chips; however systematic methodologies for quantifying these improvements lag behind hardware development cycles. We contribute an open-source microbenchmark suite that offers practical insights into optimizing workloads to fully utilize the rich feature sets of the modern GPU architecture. This work aims to enable application developers make informed architectural decisions and guide future GPU design directions.
Our work studies Blackwell GPUs, compares them to H200 generation with regards to the memory subsystem, tensor core pipeline and floating-point precisions (FP32, FP16, FP8, FP6, FP4). Our systematic evaluation of dense/sparse GEMM, transformer inference, and training workloads demonstrate that B200’s tensor core enhancements achieves 1.56× higher mixed-precision throughput and 42% better energy efficiency than H200. Our memory analysis reveals 58% reduction in memory access latency in cache-misses, fundamentally changing optimal algorithm design strategies.
@@ I Introduction
Artificial Intelligence (AI) and high‑performance computing (HPC) have evolved into data‑intensive disciplines that continuously challenge hardware efficiency, scalability, and precision. Large language models (LLMs) now exceed hundreds of billions of parameters and process context windows spanning millions of tokens [ding2024longrope, kevian2024capabilitieslargelanguagemodels], alongside multi‑physics and climate simulations that demand teraflops of sustained performance, thus shifting GPU design to enable both massive parallel and architectural adaptability. At these scales, modern accelerators must balance several demands: maintaining arithmetic throughput for dense tensor workloads, minimizing on‑chip and off‑chip memory latency, while offering hardware primitives that effectively support mixed‑precision computation.
The growing demands have exposed several limitations of current GPU architectures, particularly within their memory hierarchies, precision flexibility, and latency‑sensitive task scheduling. As a result, sustained architectural innovation in accelerators has become essential for advancing both throughput‑optimized training and time‑critical inference workloads. One such architecture that is designed to address some of these challenges is NVIDIA’s Blackwell architecture [NVIDIA2024_Blackwell] that showcases a major generational evolution.
As the direct successor to the Hopper generation, the Blackwell architecture extends NVIDIA’s GPU design in several modifications across the compute pipeline, memory hierarchy, and tensor processing subsystems. Blackwell introduces 5th‑generation Tensor Cores capable of FP4 and FP6 precision execution, offering trade‑offs between accuracy and performance for large‑scale training. In addition, the introduction of the Tensor Memory (TMEM) subsystem as a dedicated on‑chip memory for tensor data movement reduces reliance on shared memory (SMEM) and per SM register files (RF) during matrix‑intensive operations. Next, NVIDIA included a hardware decompression engine (DE) and redesigned the instruction pipeline for access to compressed model weights. Beyond raw compute enhancements, Blackwell also has a revised thread and CTA scheduling model to utilize inter‑SM communication and memory concurrency. With so many changes introduced, intended to address the escalating demands of AI, gaming and scientific computing, an analysis of the microarchitecture and new instructions is necessary, which will provide application developers and scientists to achieve the highest performance possible for modern and future GPUs.
@@
Analyze 5th-gen Tensor Core execution via the new tcgen05 PTX instructions to study performance implications
@@ II Related Work
Understanding GPU performance has long been a critical focus in HPC research. Over the years, several studies have used microbenchmarks and other methodologies to dissect architectural layers and analyze GPU microarchitectures in fine-grained detail. Early studies on Tesla and Fermi focused on memory and cache behavior [Wong2010, Subramoniapillai2012], while later work dissecting Kepler, Pascal [Zhang2017], and Maxwell [Mei2012] examined warp scheduling and instruction latency. With Turing through Hopper [Fasi2021, jia2019, tan2011, Markidis2018, Martineau2018, Raihan2019, Yan2020, luo2025], research shifted to mixed-precision and tensor core performance, introducing benchmarks for mma instructions, tile sizes, and data layouts. Recent efforts also explore instruction-level parallelism[Sun2022], and pipeline dynamics under high register pressure.
Beyond microbenchmarking, researchers built frameworks to characterize GPU performance. Application profiling[5336219] gathers runtime metrics but faces overhead and limited architectural visibility. Roofline models [leinhauser2021metricsdesigninstructionroofline] offer throughput vs. intensity plots, yet oversimplify bottlenecks and miss dynamic memory behaviors. Cache stall prediction [wenhao2012] estimates pipeline delays from access patterns but fails to capture modern GPU complexities like cache bypassing, warp scheduling, and memory coalescing.
Analytical models like Accel-Sim [accelsim2020] and GCoM [gcom2022], built on Hong and Kim’s work [Hong2009], offer useful GPU performance insights but neither of them model Blackwell-specific features like TMEM or the DE, as the detailed architectural information required for accurate simulation remain unknown.
@@ III-A Blackwell Architecture
The B200 GPU signifies a decisive progression in architectural philosophy. Previously, GPU generations from Tesla to Hopper focused on maximizing floating-point operations per second (FLOPS) for large-scale model training. In contrast, Blackwell emphasizes post-training and inference efficiency, adopting transformational changes in both memory and compute organization. One B200 GPU includes a dual-die configuration [NVIDIA2024_Blackwell] where two GPU dies comprise 208 billion transistors, feature 148 SMs spread across eight GPCs, provide four L2 cache partitions (double those in Hopper), and include eight HBM3e memory stacks. Though physically partitioned, both dies are unified by the NVIDIA High-Bandwidth Interface (NV-HBI) providing a coherent and single device to software, with unified 192 GB HBM3e memory space.
Within each SM, Blackwell introduces fifth-generation Tensor Cores that break from the warp-synchronous paradigm characterizing earlier architectures (Volta, Ampere, Hopper). Previous generations enforced that all 32 threads within a warp synchronize before executing matrix multiply–accumulate (MMA) operations via the mma.sync or wgmma instructions. This lock-step model reduced scheduling flexibility and created idle cycles, especially for dependency chains of varying lengths.
@@
Further broadening the functionality, Blackwell’s Tensor Cores provide native support for convolution operators with weight-stationary dataflows that use a collector buffer to cache and reuse matrix B (weight tensor) operands. Hence, optimizing for convolutional kernels that benefit from operand locality. Blackwell also addresses growing model and data sizes by introducing a hardware-based Decompression Engine (DE) to offload decompression tasks from general-purpose SMs. This subsystem supports various algorithms, more on that in Section V, enabling model weights and large database tables to be stored in compressed form within HBM3e and decompressed transparently during memory access [NVIDIA2024_Blackwell].
@@ IV PTX-Microbenchmark Methodology
We employ a microbenchmarking approach based on NVIDIA’s Parallel Thread Execution (PTX) to characterize Blackwell’s microarchitectural features. While prior GPU characterization studies [Wong2010, Mei2012, Yan2020] established foundational PTX-level benchmarking principles, we extend these methodologies with novel techniques specifically designed to dissect Blackwell’s previously uncharacterized components, including the 5th-generation Tensor Cores’s FP4/FP6 precision modes, the DE, and the revised cache hierarchy.
Our approach leverages PTX to provide explicit control over registers and memory operations specific to architectures. PTX code is compiled into Streaming Assembler (SASS) instructions or machine code executed by the GPU. We ensure our microbenchmarks accurately isolate and measure specific microarchitectural behaviors by documenting the PTX-to-SASS translation and validating against observed performance.
@@ IV-A1 Tensor Memory (TMEM)
Unlike previous architectures where MMA operations relied exclusively on SMEM, DSMEM, and RFs, Blackwell introduces TMEM as a dedicated on-chip memory specifically for tensor operations. This necessitates new data movement paradigms and presents unexplored opportunities for performance optimization. Understanding TMEM’s performance characteristics is critical for several reasons. First, traditional data movement instructions (including wmma.load, ldmatrix, ld.shared, and cp.async) which cannot interface with TMEM. Thus developers are required to adopt entirely new instruction sequences (tcgen05.ld, tcgen05.st, tcgen05.cp). Second, the performance implications of this new memory tier remain uncharacterized, leaving application developers without guidance on when and how to leverage TMEM effectively.
@@
Establish performance baselines by comparing memory access latencies between traditional shared memory and TMEM using pointer-chase benchmarks. This microbenchmark isolates latency effects by creating dependent memory accesses that prevent pipeline overlap, revealing the fundamental access costs of each memory tier.
@@ IV-A2 Decompression Engine Characterization
To systematically characterize the B200’s hardware DE, we develop a custom microbenchmarking suite targeting seven compression formats (LZ4, Snappy, Zstandard, GZIP, Cascaded, Bitcomp, ANS) across controlled test conditions. We measure end-to-end decompression throughput for 100MB datasets using each supported format. Input throughput is calculated as compressed data read rate from GPU memory; output throughput measures decompressed data generation rate. Latency captures complete decompression time including memory transfers. Each measurement averages 1000 iterations after 100-iteration warmup to ensure thermal and cache stability. We generate synthetic datasets with varying entropy: random data (incompressible, 1.00x ratio), mixed alphanumeric (1.98x), repetitive patterns (15.02x), and zero-filled buffers (245.45x).
All experiments use LZ4 format to isolate compression-ratio effects from format-specific optimizations. Data is pre-compressed on CPU using reference implementations, then loaded to GPU memory for DE processing. This design separates compression algorithm effects from hardware decompression characteristics.
@@ IV-A3 Tensor Core Characterization
We conduct latency and throughput measurements across varying instruction types, matrix tile shapes, and operand layouts to characterize execution pipeline behavior. Power efficiency analysis compares compute throughput against board-level power consumption to identify energy-optimal operating points for different precision modes and tile configurations.
To test the single thread instruction improvement instruction we issue multiple mma tile shapes all with FP16 inputs and outputs. We first isolate the instructions and show single instruction latency (SI-LAT). As well as compare against various levels of warps, which will account for wgmma only being issued at warp-group level which is 4 asynchronous warps per group. As well as using the TMEM as minimal as allowed, which means only for the accumulator, denoted as SS.
@@ IV-A4 Extended Precision Characterization
Unlike prior work focusing on FP8, FP16, and INT8 tensor operations [Fasi2021, Markidis2018], we develop the first systematic benchmarks for Blackwell’s FP4 and FP6 MMA instructions using the tcgen05 PTX opcode with e2m1 (FP4), e3m2 (FP6), and e2m3 (FP6) encoding formats. Our key methodological contribution is a dependency-chain approach that isolates true instruction latency for these ultra-low-precision operations.
@@ IV-A5 Workflow Benchmarks
First, we select the Mistral model family [jiang2023mistral7b] as an LLM for several reasons: (1) Mistral-7B provides a representative dense decoder architecture with performance comparable to larger models, (2) Mixtral-8x7B’s Mixture-of-Experts (MoE) architecture exercises different dataflow patterns that stress Blackwell’s memory hierarchy, and (3) the Mistral family’s public availability enables reproducibility. The architectural diversity spanning dense (Mistral-7B) to sparse MoE (Mixtral-8x7B, Mixtral-8x22B) provides comprehensive coverage of modern LLM deployment scenarios.
Next, we develop custom matrix multiplication kernels using FP64 to measure the realistic performance for scientific workloads. In addition, we run STREAM Triad [McCalpin1995] for memory bandwidth and SpMV tests with real-world data to benchmark the DE. Finally, we measure end-to-end training performance using mixed-precision training with ResNet-50 [he2015deepresiduallearningimage], and GPT-1.3B [gao2020pile].
@@ V Memory Subsystem
In this Section, we present a comparative evaluation of the memory subsystems, including TMEM and DE, through microbenchmarking methodologies that measure latency, saturation behavior, and sensitivity to access patterns. These new characteristics could fundamentally alter data movement patterns and memory bandwidth utilization compared to prior architectures.
@@ V-A Tensor Memory (TMEM)
The TMEM is a dedicated 256KB on-chip memory per SM designed solely for Tensor Core operations. Structured as a 2D array of 512 columns by 128 lanes of 32-bit cells, TMEM uses a lane-column addressing scheme [NVIDIA2024_Blackwell]. The TMEM separates tensor core storage from registers, enabling intermediate matrix results to persist across warp groups and while reducing reliance on global or shared memory.
Our latency characterization reveals that TMEM achieves 420 clock cycles for end-to-end memory access in cache-miss scenarios, representing a 58% reduction compared to Hopper’s 1000-cycle global memory latency. This improvement stems from TMEM’s dedicated arbitration logic that bypasses the L2 cache partitioning contention inherent in traditional memory hierarchies. More critically, TMEM provides 16 TB/s read bandwidth and 8 TB/s write bandwidth per SM, and this bandwidth operates additively with L1/SMEM bandwidth rather than competing for the same resources. In our microbenchmarks using back-to-back MMA operations on FP8 data, TMEM sustained 8 TB/s memory bandwidth matching HBM3e peak performance, achieving a 2.1× improvement over conventional ld.global paths which plateau at 3.8 TB/s due to L1/L2 traversal overhead.
The transition to TMEM necessitates entirely new instruction sequences, as traditional data movement instructions cannot interface with this memory tier. On Hopper, the standard pipeline for tensor operations relied on cp.async.bulk.tensor.2d for asynchronous 2D tile copies from global to shared memory, followed by ldmatrix or wmma.load to stage operands into registers before MMA execution. These instructions formed a well-understood pipeline: global memory copy engines populate shared memory tiles, barrier synchronization ensures data availability, and explicit load instructions would transfer operands to the register file.
@@
On Blackwell, the tcgen05 instruction family replaces this entire sequence. The tcgen05.cp instruction handles asynchronous tensor data transfers into or out of TMEM. The tcgen05.ld and tcgen05.st instructions provide specialized load/store operations between TMEM and registers or shared memory, enabling fine-grained control over data placement. Importantly, while Hopper required A and B operand matrices to traverse SMEM before consumption by tensor cores, tcgen05 allows MMA instructions to read operands from SMEM or TMEM. After which, tcgen05 will write accumulator results directly to TMEM, creating an asymmetric but more efficient data flow.
@@
Our instruction-level analysis across varying operand sizes and access strides reveals critical performance characteristics. TMEM achieves optimal efficiency at 64×64 element tiles (4KB for FP8 precision), aligning with the 256KB SMEM capacity per SM and fully utilizing the 1024-bit memory interface width. This represents a significant departure from Hopper’s 32×32 optimal tile size, requiring algorithmic adjustments for kernels transitioning to Blackwell. Tiles smaller than 32×32 elements underutilize the wide memory interface, achieving only 45% of peak bandwidth, while tiles larger than 128×128 elements trigger multi-phase transfers that introduce pipeline stalls and reduce effective throughput by 30%. These bandwidth saturation points directly inform optimal kernel design: matrix multiplication kernels should decompose computations into 64×64 tiles to maximize TMEM utilization, and chained operations such as those in transformer attention mechanisms ( Q×KT followed by softmax and value multiplication) should maintain intermediate results in TMEM to exploit the 16 TB/s read bandwidth for subsequent operations.
Traditional Hopper operations exhibits a serial dependency chain: global memory fetch, L2 traversal, shared memory write, barrier wait, register load, and finally MMA execution. Each stage introduced both latency and bandwidth competition. Blackwell’s TMEM operates with independent address generation units that pre-fetch tensor tiles directly into staging buffers, enabling the tcgen05.mma instructions to overlap data movement with computation. While a warp group executes tensor operations on one TMEM tile, the copy engines asynchronously populate the next tile, achieving near-perfect double buffering. This pipeline efficiency becomes particularly evident in workloads with producer-consumer relationships between MMA operations.
@@
For chained matrix multiplications where , keeping the intermediate result in TMEM eliminates an estimated 12 TB of data movement per second on a fully utilized SM compared to Hopper’s approach of writing back to global memory.
Power efficiency measurements reveal nuanced trade-offs in TMEM utilization. For kernels that stage Matrix-D accumulators in TMEM versus traditional shared memory, we observe a 15% reduction in board-level power consumption at equivalent compute throughput for large matrix dimensions (2048×2048). This efficiency gain results from reduced L2 cache thrashing and lower DRAM traffic as intermediate results remain on-chip. However, for smaller problem sizes where the entire working set fits in L1 cache, forcing TMEM allocation introduces marginal overhead from the additional copy operations, resulting in a 3-5% power increase.
These measurements establish clear guidelines: TMEM should be prioritized for multi-stage tensor pipelines with large working sets, while traditional memory hierarchies remain optimal for single-shot operations on small matrices.

### v2

@@ Abstract
As GPU architectures rapidly evolve to meet the overcoming demands of exascale computing and machine learning, the performance implications of architectural innovations remain poorly understood across diverse workloads. NVIDIA’s Blackwell (B200) generation introduce significant architectural advances including the 5th-generation tensor cores, tensor memory (TMEM), decompression engine (DE), and dual chips; however systematic methodologies for quantifying these improvements lag behind hardware development cycles. We contribute an open-source microbenchmark suite that offers practical insights into optimizing workloads to fully utilize the rich feature sets of the modern GPU architecture. This work aims to enable application developers make informed architectural decisions and guide future GPU design directions.
Our work studies Blackwell GPUs, compares them to H200 generation with regards to the memory subsystem, tensor core pipeline and floating-point precisions (FP32, FP16, FP8, FP6, FP4). Our systematic evaluation of dense/sparse GEMM, transformer inference, and training workloads demonstrates that B200’s tensor core enhancements achieve 1.85 ResNet-50 and 1.55 GPT-1.3B mixed-precision training throughput with 32% better energy efficiency than H200.
@@ I Introduction
Artificial Intelligence (AI) and high‑performance computing (HPC) have evolved into data‑intensive disciplines that continuously challenge hardware efficiency, scalability, and precision. Large language models (LLMs) now exceed hundreds of billions of parameters and process context windows spanning millions of tokens [5, 13], alongside multi‑physics and climate simulations that demand teraflops of sustained performance, thus shifting GPU design to enable both massive parallel and architectural adaptability. At these scales, modern accelerators must balance several demands: maintaining arithmetic throughput for dense tensor workloads, minimizing on‑chip and off‑chip memory latency, while offering hardware primitives that effectively support mixed‑precision computation.
The growing demands have exposed several limitations of current GPU architectures, particularly within their memory hierarchies, precision flexibility, and latency‑sensitive task scheduling. As a result, sustained architectural innovation in accelerators has become essential for advancing both throughput‑optimized training and time‑critical inference workloads. One such architecture that is designed to address some of these challenges is NVIDIA’s Blackwell architecture [22] that showcases a major generational evolution.
As the direct successor to the Hopper generation, the Blackwell architecture extends NVIDIA’s GPU design in several modifications across the compute pipeline, memory hierarchy, and tensor processing subsystems. Blackwell introduces 5th‑generation tensor cores capable of FP4 and FP6 precision execution, offering trade‑offs between accuracy and performance for large‑scale training. In addition, the introduction of the Tensor Memory (TMEM) subsystem as a dedicated on‑chip memory for tensor data movement reduces reliance on shared memory (SMEM) and per SM register files (RF) during matrix‑intensive operations. Next, NVIDIA included a hardware decompression engine (DE) and redesigned the instruction pipeline for access to compressed model weights. Beyond raw compute enhancements, Blackwell also has a revised thread and Cooperative Thread Array (CTA) scheduling model to utilize inter-SM communication and memory concurrency. With so many changes introduced, intended to address the escalating demands of AI, gaming and scientific computing, an analysis of the microarchitecture and new instructions is necessary, which will provide application developers and scientists to achieve the highest performance possible for modern and future GPUs.
@@
Analyze 5th-generation tensor core execution via the new tcgen05 PTX instructions to study performance implications
@@ II Related Work
Understanding GPU performance has long been a critical focus in HPC research. Over the years, several studies have used microbenchmarks and other methodologies to dissect architectural layers and analyze GPU microarchitectures in fine-grained detail. Early studies on Tesla and Fermi focused on memory and cache behavior [29, 25], while later work dissecting Kepler, Pascal [31], and Maxwell [21] examined warp scheduling and instruction latency. With Turing through Hopper [6, 11, 27, 18, 19, 23, 30, 17], research shifted to mixed-precision and tensor core performance, introducing benchmarks for mma instructions, tile sizes, and data layouts. Recent efforts also explore instruction-level parallelism[26], and pipeline dynamics under high register pressure.
Beyond microbenchmarking, researchers built frameworks to characterize GPU performance. Application profiling[3] gathers runtime metrics but faces overhead and limited architectural visibility. Roofline models [16] offer throughput vs. intensity plots, yet oversimplify bottlenecks and miss dynamic memory behaviors. Cache stall prediction [10] estimates pipeline delays from access patterns but fails to capture modern GPU complexities like cache bypassing, warp scheduling, and memory coalescing.
Analytical models like Accel-Sim [14] and GCoM [15], built on Hong and Kim’s work [9], offer useful GPU performance insights but neither of them model Blackwell-specific features like TMEM or the DE, as the detailed architectural information required for accurate simulation remain unknown.
@@ III-A Blackwell Architecture
The B200 GPU signifies a decisive progression in architectural philosophy. Previously, GPU generations from Tesla to Hopper focused on maximizing floating-point operations per second (FLOPS) for large-scale model training. In contrast, Blackwell emphasizes post-training and inference efficiency, adopting transformational changes in both memory and compute organization. One B200 GPU includes a dual-die configuration [22] where two GPU dies comprise 208 billion transistors, feature 148 SMs spread across eight GPCs, provide four L2 cache partitions (double those in Hopper), and include eight HBM3e memory stacks. Though physically partitioned, both dies are unified by the NVIDIA High-Bandwidth Interface (NV-HBI) providing a coherent and single device to software, with unified 192 GB HBM3e memory space.
Within each SM, Blackwell introduces 5th-generation tensor cores that break from the warp-synchronous paradigm characterizing earlier architectures (Volta, Ampere, Hopper). Previous generations enforced that all 32 threads within a warp synchronize before executing matrix multiply–accumulate (MMA) operations via the mma.sync or wgmma instructions. This lock-step model reduced scheduling flexibility and created idle cycles, especially for dependency chains of varying lengths.
@@
Further broadening the functionality, Blackwell’s Tensor Cores provide native support for convolution operators with weight-stationary dataflows that use a collector buffer to cache and reuse matrix B (weight tensor) operands. Hence, optimizing for convolutional kernels that benefit from operand locality. Blackwell also addresses growing model and data sizes by introducing a hardware-based Decompression Engine (DE) to offload decompression tasks from general-purpose SMs. This subsystem supports various algorithms, more on that in Section V, enabling model weights and large database tables to be stored in compressed form within HBM3e and decompressed transparently during memory access [22].
@@ IV PTX-Microbenchmark Methodology
We employ a microbenchmarking approach based on NVIDIA’s Parallel Thread Execution (PTX) to characterize Blackwell’s microarchitectural features. While prior GPU characterization studies [29, 21, 30] established foundational PTX-level benchmarking principles, we extend these methodologies with novel techniques specifically designed to dissect Blackwell’s previously characterized components, including the 5th-generation tensor cores’s FP4/FP6 precision modes, the DE, and the revised cache hierarchy.
Our approach leverages PTX to provide explicit control over registers and memory operations specific to architectures. PTX code is compiled into Streaming Assembler (SASS) instructions. We design dependency-controlled kernels to isolate the targeted behaviors, and validate the PTX-to-SASS translation.
@@ IV-A1 Tensor Memory (TMEM)
Unlike previous architectures where MMA operations relied exclusively on SMEM, DSMEM, and RFs, Blackwell introduces TMEM as a dedicated on-chip memory specifically for tensor operations. This necessitates new data movement paradigms and presents unexplored opportunities for performance optimization. Understanding TMEM’s performance characteristics is critical for several reasons. First, traditional data movement instructions (including wmma.load, ldmatrix, ld.shared, and cp.async) which cannot interface with TMEM. Thus developers are required to adopt entirely new instruction sequences (tcgen05.ld, tcgen05.st, tcgen05.cp). Second, the performance implications of this new memory tier remain characterized, leaving application developers without guidance on when and how to leverage TMEM effectively.
@@
Establish performance baselines by comparing memory access latencies between traditional shared memory and TMEM, using pointer-chase benchmarks [28]. To isolates access latency of each memory tier, we use dependent pointer-chase loads by creating dependent memory accesses that prevent pipeline overlap.
@@ IV-A2 Decompression Engine Characterization
To systematically characterize the B200’s hardware DE, we develop a custom microbenchmarking suite targeting seven compression formats (LZ4, Snappy, Zstandard, GZIP, Cascaded, Bitcomp, ANS) across controlled test conditions. We measure end-to-end decompression throughput for 100MB datasets using each supported format. Input throughput is calculated as compressed data read rate from GPU memory; output throughput measures decompressed data generation rate. Latency captures complete decompression time including memory transfers. To isolate DE behavior from compression overhead, all datasets are pre-compressed on the CPU and the benchmark measures only device-side decompression. Each measurement averages 1000 iterations after 100-iteration warmup to ensure thermal and cache stability. We generate synthetic datasets with varying entropy: random data (incompressible, 1.00x ratio), mixed alphanumeric (1.98x), repetitive patterns (15.02x), and zero-filled buffers (245.45x).
@@ IV-A3 Tensor Core Characterization
We conduct latency and throughput measurements across varying instruction types, matrix tile shapes, and operand layouts to characterize execution pipeline behavior. To isolate instruction latency, we use dependency chains carried through the accumulator so each MMA depends on the prior result, preventing overlap; throughput is measured with independent MMAs to saturate the tensor-core pipeline. Power efficiency analysis compares compute throughput against board-level power consumption to identify energy-optimal operating points for different precision modes and tile configurations.
@@ IV-A4 Extended Precision Characterization
Unlike prior work focusing on FP8, FP16, and INT8 tensor operations [6, 18], we develop the first systematic benchmarks for Blackwell’s FP4 and FP6 MMA instructions using the tcgen05 PTX opcode with e2m1 (FP4), e3m2 (FP6), and e2m3 (FP6) encoding formats. We use dependency chains carried through the destination operand to prevent independent issues and expose true dependent latency for each FP4/FP6 instruction variant.
@@ IV-A5 Workflow Benchmarks
First, we select the Mistral model family [12] as an LLM for several reasons: (1) Mistral-7B provides a representative dense decoder architecture with performance comparable to larger models, (2) Mixtral-8x7B’s Mixture-of-Experts (MoE) architecture exercises different dataflow patterns that stress Blackwell’s memory hierarchy, and (3) the Mistral family’s public availability enables reproducibility. The architectural diversity spanning dense (Mistral-7B) to sparse MoE (Mixtral-8x7B, Mixtral-8x22B) provides comprehensive coverage of modern LLM deployment scenarios.
Next, we develop custom matrix multiplication kernels using FP64 to measure the realistic performance for scientific workloads. In addition, we run STREAM Triad [20] for memory bandwidth and SpMV tests with real-world data to benchmark the DE. Finally, we measure end-to-end training performance using mixed-precision training with ResNet-50 [8], and GPT-1.3B [7].
@@ V Memory Subsystem
In this Section, we present a comparative evaluation of the memory subsystems, including TMEM and DE, through microbenchmarking methodologies that measure latency, saturation behavior, and sensitivity to access patterns. These new characteristics can significantly change data movement patterns and effective bandwidth utilization for workloads that (i) stage tensor operands/accumulators through TMEM and (ii) stream compressed inputs through the DE, compared to prior architectures.
@@ V-A Tensor Memory (TMEM)
The TMEM is a dedicated 256KB on-chip memory per SM designed solely for Tensor Core operations. Structured as a 2D array of 512 columns by 128 lanes of 32-bit cells, TMEM uses a lane-column addressing scheme [22]. The TMEM separates tensor core storage from registers, enabling intermediate matrix results to persist across warp groups and reducing reliance on registers or shared memory.
@@
The transition to TMEM necessitates entirely new instruction sequences, as traditional data movement instructions cannot interface with this memory tier. On Hopper, the standard pipeline used cp.async.bulk.tensor.2d (or TMA) for global-to-shared tile copies, then ldmatrix or wmma.load to stage operands; the wgmma instruction accepts operand A from registers or shared memory and operand B from shared memory (B is not staged into registers before MMA). On Blackwell, the tcgen05 instruction family replaces this entire sequence. The tcgen05.cp instruction handles asynchronous tensor data transfers into or out of TMEM. The tcgen05.ld and tcgen05.st instructions provide specialized load/store operations between TMEM and registers or shared memory, enabling fine-grained control over data placement. Importantly, while Hopper required A and B operand matrices to traverse SMEM before consumption by tensor cores, tcgen05 allows MMA instructions to read operands from SMEM or TMEM. After which, tcgen05 will write accumulator results directly to TMEM, creating an asymmetric but more efficient data flow.
@@
Our instruction-level analysis across varying operand sizes and access strides reveals critical performance characteristics. TMEM achieves optimal efficiency at 6464 element tiles (4KB for FP8 precision), aligning with the 256KB TMEM capacity per SM and fully utilizing the 1024-bit memory interface width. (PTX wgmma matrix shapes require M64, e.g. m64n8k16 through m64n256k16.) Tiles smaller than 3232 elements underutilize the wide memory interface in our TMEM microbenchmarks, while tiles larger than 128128 elements can trigger multi-phase transfers. These bandwidth saturation points directly inform optimal kernel design: matrix multiplication kernels should decompose computations into 64×64 tiles to maximize TMEM utilization, and chained operations such as those in transformer attention mechanisms ( followed by softmax and value multiplication) should maintain intermediate results in TMEM to exploit the 16 TB/s read bandwidth for subsequent operations.
Traditional Hopper operations exhibit a serial dependency chain: global memory fetch (e.g. TMA or cp.async.bulk), shared memory write, barrier wait, then wgmma consuming operands from shared (and optionally registers). On Blackwell, TMEM is software-managed via tcgen05.alloc, tcgen05.cp, tcgen05.ld/tcgen05.st; software uses tcgen05.cp to copy data into TMEM from shared memory, and tcgen05.mma can then overlap with subsequent tcgen05.cp for double-buffered pipelines.
@@
For chained matrix multiplications where , keeping the intermediate result in TMEM eliminates a substantial amount of off-chip traffic relative to a Hopper-style design that writes intermediates back to global memory. Under bandwidth and traffic assumptions, this corresponds to an estimated of data movement per SM that can be avoided on a fully utilized Blackwell SM.